# Token-Importance Guided Direct Preference Optimization

### v1

@@ 4.3 TI-DPO Objective and Theoretical Analysis
DPO defines the reward model as the log-likelihood ratio between $\pi_{\theta}$ and $\pi_{\text{ref}}$, eliminating the need for an explicit reward function. However, TI-DPO extends this by incorporating triplet loss, which explicitly guides outputs $y$ to align with preferred responses $y_{\text{w}}$ and diverge from non-preferred $y_{\text{l}}$. The dataset $D=\{(x,y_{\text{w}},y_{\text{l}})\}$ enables learning to pull responses toward high-quality standards and reject low-quality alternatives. Building on the triplet loss definition in Eq.(5), we obtain the triplet objective with policy model as follows:
$\mathcal{L}_{\text{triplet}}=$ $\mathbb{E}_{\left(x,y_{\text{w}},y_{\text{l}}\right)\sim\mathcal{D}}\Big{[}\max(0,\ \underbrace{\sum_{t=1}^{T_{\text{w}}}\Big{\|}\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}-\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\Big{\|}^{2}_{2}}_{\text{Bring }y\text{ closer to }y_{\text{w}}}$ (14)
$-\underbrace{\sum_{t=1}^{T_{\text{l}}}\Big{\|}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}-\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}\Big{\|}^{2}_{2}}_{\text{Push }y\text{ away }y_{\text{l}}}+\alpha)\Big{]}_{+}.$
@@
where $\gamma$ is a hyperparameter. This combined loss guides the policy $\pi_{\theta}$ to learn to assign similar scores to $y_{\text{w}}$, while emphasizing the importance of tokens in $y_{\text{w}}$ through the weighted reward function. In Appendix A.3, we have given the proof of gradient $\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}$, which is used to update $\theta$ during training. Denote $\Delta r_{\text{global}}=\log\frac{\pi_{\theta}\left(y_{\text{w}}|x\right)}{\pi_{\text{ref}}\left(y_{\text{w}}|x\right)}-\log\frac{\pi_{\theta}\left(y_{\text{l}}|x\right)}{\pi_{\text{ref}}\left(y_{\text{l}}|x\right)}$ from Eq.(4). For simplicity, we abbreviate $\Delta r_{\text{token}}(x,y_{\text{w}},y_{\text{l}},w^{\text{w}}_{t},w^{\text{l}}_{t})$ as $\Delta r_{\text{token}}$. To show the superiority of our TI-DPO loss compared to DPO loss, we first introduce the following lemma:
Lemma 4.1
There is $\alpha>1$ s.t. $\forall(x,y_{\text{w}},y_{\text{l}})\sim D$, it is satisfied that:
$\mathbb{E}\left[\Delta r_{\text{token}}\right]\geq\alpha\cdot\mathbb{E}\left[\Delta r_{\text{global}}\right].$ (16)
The proof of Lemma 4.1 is detailed in Appendix A.1. According to Eq.(13), TI-DPO dynamically allocates token importance weights through gradient attribution, enabling the model to focus on the generation outputs with more importance weights on preference alignment. The following theorem strictly proves the theoretical advantages of this improvement at the loss function level. Under the condition of Lemma 4.1, the total loss of TI-DPO will be significantly lower than the original DPO loss:
Theorem 4.2 (Tighter Loss Bound)
If $\exists\alpha>1$ s.t. $\forall(x,y_{\text{w}},y_{\text{l}})\sim D$ satisfying $\mathbb{E}\left[\Delta r_{\text{token}}\right]\geq\alpha\cdot\mathbb{E}\left[\Delta r_{\text{global}}\right]$, then TI-DPO’s total loss is strictly lower than DPO’s, i.e.,
$\mathcal{L}_{\text{TI-DPO}}\leq\mathcal{L}_{\text{DPO}}-\beta\Delta_{\text{triplet}},$ (17)
where $\Delta_{\text{triplet}}$ is a positive coefficient related to triplet loss and $\alpha$.
The proof is shown in Appendix A.2. Theorem 4.2 indicates that the total loss of TI-DPO is strictly less than the original DPO loss by an additional positive term determined jointly by the token-level weight advantage and the triplet coefficient. The strict superiority of $\mathcal{L}_{\text{TI-DPO}}$ provides a mathematical guarantee for the more stable training process of TI-DPO in practical tasks, verifying its essential improvement compared to the traditional DPO from a theoretical perspective. The implementation of TI-DPO is shown in Algorithm 1 as follows.
@@
1:Input: Dataset $\mathcal{D}=\{(x,y_{\text{w}},y_{\text{l}})\}$, hyperparameter $\beta$ and $\alpha$, the length of preferred responses $T_{\text{w}}$, the length of non-preferred responses $T_{\text{l}}$, reference model $\pi_{\text{ref}}$, policy model $\pi_{\theta}$.
@@
5:     Compute policy model $\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})$ and $\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})$.
6:     Compute reference model $\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})$ and $\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})$.
7:     Calculate importance weights for preferred responses:
8:     $w_{t}^{\text{w}}=\text{Normalize}(\|\frac{1}{T_{\text{w}}}\sum_{t=1}^{T_{\text{w}}}\nabla_{e_{t}}\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\|_{1}).$
9:     Calculate importance weights for non-preferred responses:
10:     $w_{t}^{\text{l}}=\text{Normalize}(\|\frac{1}{T_{\text{l}}}\sum_{t=1}^{T_{\text{l}}}\nabla_{e_{t}}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}\|_{1}).$
11:     Compute weighted DPO log-ratio:
@@
12:     $\Delta r_{\text{token}}=\sum_{t=1}^{T_{\text{w}}}w^{\text{w}}_{t}\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}-\sum_{t=1}^{T_{\text{l}}}w^{\text{l}}_{t}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}.$
13:     Calculate the output $y^{t}$: $y^{t}\sim\pi_{\theta}(y^{t-1}|x,y^{<t-1})$.
14:     Compute triplet log-ratio:
15:     $\Delta r_{\text{triplet}}=\sum_{t=1}^{T_{\text{w}}}\Big{\|}\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}-\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}\Big{\|}^{2}_{2}-\sum_{t=1}^{T_{\text{l}}}\Big{\|}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}-\log\frac{\pi_{\theta}(y^{t}|x,y^{<t})}{\pi_{\text{ref}}(y^{t}|x,y^{<t})}\Big{\|}^{2}_{2}$.
@@
16:     Compute weighted DPO loss: $\mathcal{L}_{\text{DPO-w}}=-\log\sigma(\beta\Delta r_{\text{token}})$.
@@
17:     Compute triplet loss: $\mathcal{L}_{\text{triplet}}=\max\left(0,\Delta r_{\text{triplet}}+\alpha\right)$.
18:     Aggregate losses: $\mathcal{L}_{\text{TI-DPO}}\leftarrow\mathcal{L}_{\text{DPO-w}}+\gamma\mathcal{L}_{\text{triplet}}$.
19:     Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}$.
20:end for
21:Output: $\pi_{\theta}$
5 Experiments
This section conducts extensive experiments to evaluate our proposed TI-DPO. Our experiments show the effectiveness of token-importance guided weight and analyze the performance on model alignment compared with other works.
5.1 Experimental Settings
Dataset and base settings: The following datasets are employed across various tasks to evaluate our approach. MMLU [43] is used to assess the model’s knowledge breadth, covering multiple-choice tasks in 57 disciplines. GSM8K [44] and MATH [45] focus on mathematical reasoning, including primary school math application problems and advanced math problems, respectively. MT-Bench [47] and its enhanced version Arena-Hard are judged by GPT-4, examining the conversation ability and complex instruction processing. Additionally, TruthfulQA [49] detects the authenticity of the model’s answers through adversarial questions. GPQA [50] evaluates graduate-level scientific question-answering with 448 multi-disciplinary multiple-choice questions, and IFEval [51] measures instruction-following accuracy through  500 verifiable tasks. HumanEval [52] examines the code generation ability with the pass@1 indicator.
Comparative algorithm: We compared the TI-DPO with baseline alignment methods such as SFT, DPO [14], IPO[53], KTO [33], SimPO [19], TDPO [17], CPO [16], TPO [25], and GRPO [28]. In different alignment tasks, we selected a base model that balances performance and efficiency, comprehensively considering the model parameter scale and adaptability to guarantee the fairness and efficacy of the comparative experiment. Specifically, we select three models (Llama-3.2-3B [54], Llama-3.1-8B [54], Mistral-7B-v0.3 [55]) as baselines.
5.2 Performance Comparison
(a) TruthfulQA Performance vs. Training Steps
(b) IFEval Performance vs. Training Steps
As shown in Figure 2, TI-DPO exhibits significantly better performance in Reasoning, Instruction-Following, and Reliability dimensions compared to the corresponding instruction variants with each base instruct. Here, the base instruct refers to the foundational instruction-tuned models (Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3), serving as the baseline for comparing the effectiveness of TI-DPO and other fine-tuning methods. The scores of TI-DPO in other aspects are roughly equal or slightly higher than others. Table 1 presents average scores of each fine-tuning method across three base models, clearly demonstrating our method’s advantages in general tasks and specific scenarios. The specific score comparison table under the three base model is placed in Appendix B.
@@ 5.3 Ablation Experiment
Base Instruct (Baseline) 0.634 0.777 0.266 0.280 0.515 0.768
Full Method (TI-DPO) 0.654 0.807 0.346 0.330 0.635 0.868
No Triplet Loss 0.640 0.790 0.320 0.310 0.605 0.830
Uniform Weight 0.640 0.782 0.305 0.290 0.580 0.800
Random Weight 0.637 0.778 0.280 0.285 0.550 0.780
In the ablation experiment section, we conduct an ablation analysis on the impact of the "importance guidance" mechanism and the "triplet loss" on the model performance. It focuses on tasks where the impact of the expected importance mechanism such as GSM8K [44] for mathematical reasoning and TruthfulQA [49] for reliability is significant.
In order to verify the improvement effect of various weight and triplet conditions, Table 2 shows the ablation experiment with Llama-3.2-3B-Instruct as baseline. The experimental design compares TI-DPO (full method), baseline with randomized weights, baseline with uniform importance weights but ignoring the importance differences of tokens, as well as different configurations with or without the triplet loss. It can be observed that the ordered design of weight (from Baseline $\rightarrow$ Random weight $\rightarrow$ Uniform Weight $\rightarrow$ Full Method) leads to a gradually enhanced performance in all aspects. The design of the triplet loss comprehensively improves the performance score, especially for Instruction-Following and Reliability.
5.4 Case Demo
We present a medical scenario case demo in Fig. 3 to illustrate how TI-DPO dynamically assigns token importance weights and leverages triplet loss to optimize model outputs. Given the user prompt "I have a headache, what should I do?", three responses are analyzed: preferred response (A), intermediate response (B), and non-preferred response (C). More details of this demo would be shown in Appendix B.2.
Figure 3: Case demo of responses to prompt "I have a headache, what should I do?". Left: Preferred case. Middle: Inermedia case. Right: Non-preferred case. The darker color indicates higher weight.
@@ 6 Conclusion
We introduce TI-DPO, an optimization framework that effectively bridges the alignment gap between LM and human value systems. By incorporating token-importance guidance and a triplet loss, TI-DPO effectively overcomes the limitations of traditional DPO methods, which often neglect individual token significance and are sensitive to judgment noise. Rigorous experiments across diverse benchmarks such as HumanEval, TruthfulQA, and IFEval show that TI-DPO consistently outperforms baselines like SFT, DPO, and their variants. In performance comparisons, TI-DPO achieves the highest accuracy on TruthfulQA and IFEval tasks, demonstrating its superiority in fine-grained preference alignment. And we compare TI-DPO with three base instruction models across multiple task dimensions with radar graphs. Additionally, ablation experiments confirm the necessity of its token-importance guidance and triplet loss components. Visualized weight distributions in the appendix reveal that TI-DPO focuses on semantically critical tokens, such as safety terms of medical responses that we show in the case study, while theoretical analysis proves it achieves a tighter loss bound than DPO, ensuring optimization stability. These experiments highlight the good performance of TI-DPO in improving the quality and reliability of LLM output, providing a powerful solution for preference optimization.

### v2

@@ 4.4 TI-DPO Objective and Theoretical Analysis
where $\gamma$ is a hyperparameter. In Appendix A.4, we have given the proof of gradient $\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}$, which is used to update $\theta$ during training. The implementation of TI-DPO is shown in Algorithm 1.
To show the superiority of TI-DPO, we first introduce the following lemma.
Lemma 1 (Variance Reduction).
Consider a reward signal governed by a sparse set of critical tokens, such that the subset of non-critical tokens $\mathcal{N}$ contributes only independent zero-mean noise $\epsilon_{t}$ with variance $\sigma^{2}_{\epsilon}$. Provided that the importance weights for these non-critical tokens are suppressed such that $w_{t}^{2}<1$ for all $t\in\mathcal{N}$, the variance of the TI-DPO estimator ($\sigma^{2}_{\text{TI-DPO}}$) is strictly lower than that of the standard DPO estimator ($\sigma^{2}_{\text{DPO}}$), i.e.,
$\sigma^{2}_{\text{TI-DPO}}<\sigma^{2}_{\text{DPO}}$ (14)
The proof is detailed in Appendix A.1. Lemma 1 establishes that by suppressing the weights of non-critical tokens, TI-DPO effectively filters out stochastic noise, resulting in a strictly lower variance for the reward difference estimator. The following theorem strictly proves the theoretical advantages of this improvement at the loss function level. With Lemma 1, the total loss of TI-DPO will be significantly lower than the original DPO loss. Denote $\Delta r_{\text{global}}=\log\frac{\pi_{\theta}\left(y_{\text{w}}|x\right)}{\pi_{\text{ref}}\left(y_{\text{w}}|x\right)}-\log\frac{\pi_{\theta}\left(y_{\text{l}}|x\right)}{\pi_{\text{ref}}\left(y_{\text{l}}|x\right)}$ from Eq.(3). For simplicity, we abbreviate $\Delta r_{\text{token}}(x,y_{\text{w}},y_{\text{l}},w^{\text{w}}_{t},w^{\text{l}}_{t})$ as $\Delta r_{\text{token}}$, then we have:
Theorem 2 (Tighter Loss Bound).
Assuming the preference optimization objective is a strictly convex function and the reward difference estimation is unbiased, the expected loss of TI-DPO ($\mathcal{L}_{\text{TI-DPO}}$) is strictly upper-bounded by that of DPO ($\mathcal{L}_{\text{DPO}}$) minus a term proportional to the variance reduction. Specifically:
$\mathcal{L}_{\text{TI-DPO}}\leq\mathcal{L}_{\text{DPO}}-\frac{1}{2}\kappa\Delta_{\sigma^{2}},$ (15)
where $\kappa>0$ represents the lower bound of the loss function’s local curvature, and $\Delta_{\sigma^{2}}=\sigma^{2}_{\text{DPO}}-\sigma^{2}_{\text{TI-DPO}}$ is the positive variance reduction term derived in Lemma 1.
The proof is shown in Appendix A.2. In the experiments presented in the Appendix B.5, we compared the loss function convergence processes of the TI-DPO and DPO, thereby further substantiating Theorem 2. Furthermore, we provide a theoretical justification for the superiority of the policy learned by TI-DPO. In Theorem 3, we demonstrate that TI-DPO utilizes the limited KL constraint more efficiently by concentrating probability mass on critical tokens. The proof is shown in Appendix A.3.
Theorem 3 (Superiority of Optimal Policy).
Let $\pi_{DPO}$ and $\pi_{TI-DPO}$ be the optimal policies derived from minimizing the DPO and TI-DPO objectives, respectively. Under a fixed total KL divergence constraint $K_{total}$, the expected true reward of the TI-DPO optimal policy is strictly lower-bounded by that of the DPO policy, i.e.,
$\mathbb{E}_{y\sim\pi_{TI-DPO}}[r^{*}(x,y)]\geq\mathbb{E}_{y\sim\pi_{DPO}}[r^{*}(x,y)]+\delta,$ (16)
where $\delta>0$ represents the gain derived from optimizing the decomposition of KL divergence, specifically by minimizing the divergence component on non-critical tokens.
5 Experiments
5.1 Experimental Settings
Dataset and base settings: The benchmarks we use include knowledge-based tasks (MMLU (Hendrycks et al., 2020)), mathematical reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021)), instruction-following (IFEval (Zhou et al., 2023)), and code generation (HumanEval (Chen et al., 2021)). Additionally, TruthfulQA (Lin et al., 2021) detects the authenticity of the model’s answers through adversarial questions.
Comparative algorithm: We compared the TI-DPO with baseline alignment methods such as SFT, DPO, IPO(Azar et al., 2024), KTO (Ethayarajh et al., 2024), SimPO (Meng et al., 2024), TDPO (Zeng et al., 2024), CPO (Feng et al., 2025), TPO (Saeidi et al., 2024), TIS-DPO (Liu et al., 2024a), Logic-RL (Xie et al., 2025b), cDPO (Lin et al., 2024) and GRPO (Shao et al., 2024). We select three models (Llama-3.2-3B (Grattafiori et al., 2024), Llama-3.1-8B (Grattafiori et al., 2024), Mistral-7B-v0.3 (Jiang et al., 2023)) as baselines.
@@ 5.1 Experimental Settings
1:Input: Dataset $\mathcal{D}=\{(x,y_{\text{w}},y_{\text{l}})\}$, hyperparameter $\beta,\alpha,\lambda$, reference model $\pi_{\text{ref}}$, policy model $\pi_{\theta}$.
@@
5:  Compute raw importance scores $\mathcal{I}$ via gradient attribution.
6:  Compute weights $\{w_{t}^{\text{w}}\}$ and $\{w_{t}^{l}\}$ by mixing normalized scores with a Gaussian prior $\mathcal{P}_{\text{prior}}$: $W\leftarrow\lambda\mathcal{I}_{\text{norm}}+(1-\lambda)\textperiodcentered\mathcal{P}_{\text{prior}}$.
@@
7:  Compute weighted DPO log-ratio: $\Delta r_{\text{token}}\leftarrow\sum_{t}w^{\text{w}}_{t}\log\frac{\pi_{\theta}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}{\pi_{\text{ref}}(y_{\text{w}}^{t}|x,y_{\text{w}}^{<t})}-\sum_{t}w^{\text{l}}_{t}\log\frac{\pi_{\theta}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}{\pi_{\text{ref}}(y_{\text{l}}^{t}|x,y_{\text{l}}^{<t})}.$
8:  Generate the anchor response $y^{t}$: $y^{t}\sim\pi_{\theta}(y^{t-1}|x,y^{<t-1})$.
9:  Compute triplet log-ratio $\Delta r_{\text{triplet}}$ (Eq.(10)).
@@
10:  Compute weighted DPO loss: $\mathcal{L}_{\text{DPO-w}}\leftarrow-\log\sigma(\beta\Delta r_{\text{token}})$.
@@
11:  Compute triplet loss: $\mathcal{L}_{\text{triplet}}\leftarrow\max\left(0,\Delta r_{\text{triplet}}+\alpha\right)$.
12:  Aggregate losses: $\mathcal{L}_{\text{TI-DPO}}\leftarrow\mathcal{L}_{\text{DPO-w}}+\gamma\mathcal{L}_{\text{triplet}}$.
13:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{TI-DPO}}$.
14:end for
15:Output: $\pi_{\theta}$
@@
Base Instruct (Baseline) 63.4 77.7 26.6 28.0 51.5 76.8
Full Method (TI-DPO) 65.4 80.7 34.6 33.0 63.5 86.8
No Triplet Loss 64.0 79.0 32.0 31.0 60.5 83.0
Uniform Weight 64.0 78.2 30.5 29.0 58.0 80.0
Random Weight 63.7 77.8 28.0 28.5 55.0 78.0
No Gaussian Prior 64.5 79.7 32.7 31.5 60.0 82.5
Softmax Prior 64.2 78.8 31.8 30.0 59.0 81.0
5.2 Performance Comparison
(a) TruthfulQA Performance vs. Training Steps
(b) IFEval Performance vs. Training Steps
As shown in Figure 1, TI-DPO exhibits significantly better performance in Reasoning, Instruction-Following, and Reliability dimensions compared to the corresponding instruction variants with each base instruction. Here, the base instruct refers to the foundational instruction-tuned models (Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3), serving as the baseline for comparing the effectiveness of TI-DPO and other fine-tuning methods. The scores of TI-DPO in other aspects are roughly equal or slightly higher than others. Table 1 presents average scores of each fine-tuning method across three base models, clearly demonstrating our method’s advantages in general tasks and specific scenarios. The specific score comparison table under the three base models is placed in Appendix C.
5.3 Ablation Experiment
To verify the distinct contributions of the importance guidance and triplet loss, Table 2 presents an ablation study with Llama-3.2-3B-Instruct. The results validate the effectiveness of our design choices: compared to the Random Weight and Uniform Weight settings, the Full Method achieves the highest scores across all six dimensions. Specifically, the importance of the Triplet Loss is evidenced by the drop in Math (80.7 $\rightarrow$ 79.0) and Code (33.0 $\rightarrow$ 31.0) scores when it is removed. Similarly, ablating the Gaussian Prior leads to a notable decline in Reliability (86.8 $\rightarrow$ 82.5).
5.4 Additional Experiments
Case Study: A case study on a medical query (see Appendix C.1 for details) demonstrates that, given the user prompt “I have a headache, what should I do?”, TI-DPO effectively assigns higher importance to safety-critical tokens (e.g., “medical attention”, “promptly”) in preferred responses, while penalizing risky suggestions (e.g., “painkillers”, “casually”) in non-preferred ones. Additionally, there are two other cases in Appendix C.2.
Pearson correlation coefficient: To investigate the effectiveness of our hybrid weighting mechanism, we also conducted Pearson correlation coefficient analysis, with full results and methodology presented in Appendix B.2.
Robustness and Generation Diversity: We validate the robustness and generative diversity of TI-DPO, which can be seen in Appendix B.4.
Sensitivity of Hyperparameters: We conducte sensitivity analyses for the weight-mixing parameter $\lambda$ and KL weight $\alpha$ in Appendix B.6. And we have provided the specific values of the hyperparameters in this project, which was also shown in Appendix B.6.
@@ 6 Conclusion
We introduce TI-DPO, an optimization framework that effectively bridges the alignment gap between LLMs and human value systems. By introducing a mixed weight calculated collaboratively by gradient attribution and Gaussian prior, TI-DPO effectively overcomes the limitations of traditional DPO methods at the token level and their sensitivity to noise. On this basis, the triplet loss structure provides more refined guidance for model optimization. Theorem 2 and Theorem 3 theoretically illustrate the superiority of TI-DPO over DPO. The effectiveness of TI-DPO is unequivocally demonstrated through extensive experimentation. Our method achieves a state-of-the-art average score of 62.3 across a diverse suite of benchmarks, outperforming all baseline methods. This superiority is particularly evident in complex alignment tasks: On HumanEval, TruthfulQA and IFEval, TI-DPO scores 67.0, 62.0 and 75.7 respectively, significantly surpassing strong contenders like GRPO and traditional DPO.
As for the limitations, despite its effectiveness in fine-grained alignment, TI-DPO entails a computational overhead during training and performs slightly below sequence-level baselines on holistic reasoning tasks. Future work will focus on integrating our token-importance mechanism with group-based optimization methods like GRPO to bridge this gap and further enhance reasoning capabilities. More statements can be found in Appendix D.
In summary, TI-DPO demonstrates outstanding performance in enhancing the output quality and reliability of LLMs through its innovative design.